OpenStudy (anonymous):

y  =  -3x + 5 y  =  x2 + 2x + 9

OpenStudy (anonymous):

what do you want with this? what are you asking?

OpenStudy (anonymous):

y1 = -3x + 5 y2 = x^2 + 2x + 9. Solving means finding the coordinates (x and y) of the 2 intersections A and B of the line y1 and the parabola y2. To solve this system, make y1 = y2 and solve for x x^2 + 2x + 9 = -3x + 5 x^2 + 5x + 4 = 0. Since a - b + c = 0, one real root is (xA = -1) and the other is (xB = -c/a = -4) (TIP 2 in solving quadratic equations with shortcut). To complete find all the coordinates of the 2 intercepts. yA = -3(xA) + 5 = 3 + 5 = ...., and yB = -3(xB) + 5 = ....